OpenStudy (anonymous):

Factor completely: 45c7d3 – 30cd9 + 15cd2

OpenStudy (anonymous):

CAN YOU TAKE ME THROUGH THE STEPS

OpenStudy (anonymous):

15cd2(3c6d-2d7+d)

OpenStudy (anonymous):

ok/

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

np

OpenStudy (anonymous):

do u understand how to do it?

OpenStudy (anonymous):

\[15 c d^2 \left(3 c^6 d-2 d^7+1\right) \]

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

\[15 c d^2 \left(3 c^6 d-2 d^7+1\right)=45 c^7 d^3-30 c d^9+15 c d^2 \]